OpenStudy (anonymous):

Use Newton's method with the specified initial approximation x1 to find x3, the third approximation to the root of the given equation. (Round your answer to four decimal places.) x5 − x − 7 = 0, x1 = 1

OpenStudy (amistre64):

define the tangent line equation at the given point; and find its x intercept. use that x intercepts as a new guess for X, and repeat

OpenStudy (amistre64):

tangent y = f'(a)(x-a) + f(a) 0 = f'(a)(x-a) + f(a) 0 - f(a) = f'(a)(x-a) - f(a)/ f'(a) = x-a a - f(a)/ f'(a) = x

OpenStudy (amistre64):

\[X_2=1-\frac{1^5-1-7}{5(1)-1}\] \[X_3=X_2-\frac{(X_2)^5-(X_2)-7}{5(X_2)-1}\]

OpenStudy (amistre64):

forgot the ^4 parts

OpenStudy (anonymous):

x^4 + x - 4 = 0 ======> f (x) = x^4 + x - 4 f '(x) = 4x^3 + 1 lets start with X1 = 1 Xn+1 = Xn - [ f(x) / f '(x) ] Xn+1 = Xn - [ (X^4 + X - 4) / (4X^3 + 1) ] ======> X1 = 1 X2 = 1 - [ (1^4 + 1 - 4) / (4 * 1^3 + 1) ] ======> X1 = 1.4 X3 = 1.4 - [ ( (1.4)^4 + 1.4 - 4) / (4 * (1.4)^3 + 1) ] ======> X1 = 1.2963259853039 X4 = 1.2963259853039 - [ ( (1.2963259853039)^4 + 1.2963259853039 - 4) / (4 * (1.2963259853039)^3 + 1) ] ======> X1 = 1.2839439534061 X4 = 1.2839439534061 - [ ( (1.2839439534061)^4 + 1.2839439534061 - 4) / (4 * (1.2839439534061)^3 + 1) ] ======> X1 = 1.2837816933800 X4 = 1.2837816933800 - [ ( (1.2837816933800)^4 + 1.2837816933800 - 4) / (4 * (1.2837816933800)^3 + 1) ] ======> X1 = 1.2837816658635 THE root is 1.283782 ========= 2) x^3 - x^2 - 1 = 0 ======> f (x) = x^3 - x^2 - 1 f '(x) = 3x^2 - 2x lets start with X1 = 1 Xn+1 = Xn - [ f(x) / f '(x) ] Xn+1 = Xn - [ (X^3 - X^2 - 1) / (3X^2 - 2X) ] ======> X1 = 1 X2 = 1 - [ (1^3 - 1^2 - 1) / (3 * (1)^2 - 2 * 1) ] ======> X2 = 2 X3 = 2 - [ (2^3 - 2^2 - 1) / (3 * (2)^2 - 2 * 2) ] ======> X3 = 1.625 X4 = 1.625 - [ ( (1.625)^3 - (1.625)^2 - 1) / (3 * (1.625)^2 - 2 * 1.625) ] ======> X4 = 1.485785953177 X5 = 1.485785953177 - [ ( (1.485785953177)^3 - (1.485785953177)^2 - 1) / (3 * (1.485785953177)^2 - 2 * 1.485785953177) ] ======> X5 = 1.46595591974 X6 = 1.46595591974 - [ ( (1.46595591974)^3 - (1.46595591974)^2 - 1) / (3 * (1.46595591974)^2 - 2 * 1.46595591974) ] ======> X6 = 1.4655713749 X7 = 1.4655713749 - [ ( (1.4655713749)^3 - (1.4655713749)^2 - 1) / (3 * (1.4655713749)^2 - 2 * 1.4655713749) ] ======> X7 = 1.465571232 the root is 1.465571

OpenStudy (anonymous):

hope that helped